OpenStudy (anonymous):

12x^3y^2+9x^2y^3/3x^2y

OpenStudy (mathstudent55):

What can you factor out of the numerator?

OpenStudy (anonymous):

its suppose to be a divsion

OpenStudy (mathstudent55):

Correct, but the way to simplify it is to factor the numerator first.

OpenStudy (anonymous):

okay explain to me

OpenStudy (anonymous):

\[\frac{ 12x ^{3}y ^{2}+9x ^{2}y ^{3} }{ 3x ^{2}y }\] now what is H.C.F of 12 &9 ?

OpenStudy (anonymous):

It is 3. lowest power of x in the numerator=2 lowest power of y in the numerator=2 common in the numerator\[=3x ^{2}y ^{2}\] \[=\frac{ 3x ^{2}y ^{2}\left( 4x+3y \right) }{ 3x ^{2}y }\] now you can solve.